OpenStudy (hockeychick23):

Suppose that a population parameter is 0.3 and many large sample are taken from the population. If the sample proportions are normally distributed, with 99.7% of the sample proportions falling between 0.048 and 0.552, what is the standard deviation of the sample proportions? A. 0.094 B. 0.084 C. 0.064 D. 0.074

OpenStudy (jessiegonzales):

Margin of error = (upper limit - lower limit) / 2 = (0.582 - 0.018) / 2 = 0.282 Margin of error = Confidence coefficient coefficient for 99.7% confidence level*Standard error of p 0.282 = 3*Standard error of p Standard error of p = 0.282/3 = 0.094

OpenStudy (hockeychick23):

ohh so standard of error is the same thing as standard deviation of sample proportions?

OpenStudy (jessiegonzales):

believe so!

OpenStudy (hockeychick23):

ok thanks so much!!!

OpenStudy (jessiegonzales):

yea no problem!